An airplane maintaining a constant airspeed takes as long to go 450 mi with the wind as it does to go 375 mi against the wind. If the wind is blowing at 15 mph, what is the speed of the plane in still air?
Answer by checkley71(8403)

450/(x+15)=375/x now cross multiply
450x=375(x+15)
450x-375x=5625
75x=5625
x=5625/75
x=75 speed of the plane.
proof
450/(75+15)=375/75
450/90=5
5=5